OpenStudy (anonymous):

Replace the “?” with <, =, or > to make the statement true: 2/3 and 13/18

OpenStudy (tkhunny):

Have you considered altering 2/3? Something about a "6" comes to mind.

OpenStudy (anonymous):

Thanks for the tip.

hero (hero):

Have you considered cross multiplication to figure out <, >, or = ? \[\frac{a}{b} ? \frac{c}{d}\] \[(a)(d) ? (b)(c)\] If (a)(d) is greater then \((a)(d) > (b)(c) \) and \(\frac{a}{b} > \frac{c}{d}\) If (b)(c) is greater then, \((a)(d) < (b)(c)\) and \(\frac{a}{b} < \frac{c}{d}\) If (a)(b) = (b)(c), then \((a)(d) = (b)(c)\) and \(\frac{a}{b} = \frac{c}{d}\)